The perimeter of a triangle is 5 ft. The longest side of the triangle measures 20 in. more than the shortest side. The middle side is 3 times the measure of the shortest side. Find the lengths of the three sides in inches.
so far i have this but I'm not sure if their right
P=5 ft.?
L=S+20?
M=3s?

Answer by ankor@dixie-net.com(22740)

The perimeter of a triangle is 5 ft.
The longest side of the triangle measures 20 in. more than the shortest side.
The middle side is 3 times the measure of the shortest side.
Find the lengths of the three sides in inches.
so far i have this but I'm not sure if their right
:
P=5 ft.?
L=S+20?
M=3s?
:
This is right but notice the perimeter is in feet therefore you should have
p = 5 * 12 = 60 in.
the equation then is
(s+20) + 3s + s = 60
5s = 60 - 20
I'm sure you can take it from here